THE PERMANENT ELECTRIC DIPOLE MOMENT OF CaD ^{a}.

\underline{JINHAI CHEN}, TIMOTHY C. STEIMLE, Department of Chemistry and Biochemistry, Arizona State University, Tempe, AZ 85287.

In 2004 we determined from the analysis of the Stark effect in the A^{2}Π − X^{2}Σ^{+} (0,0) band of system the magnitude of the permanent electric dipole moment, |μ|, of CaH to be 2.94(16) D and 2.372(12)D for the X^{2}Σ^{2} (v=0) and A^{2}Π(v=0) states,respectively^{b}. The numbers in parenthesis are 1σ error estimates. Recently, two groups have performed high-level, all-electron, ab initio calculation and predicted μ values of 2.623 D^{c} and 2.617 D^{d} for the X^{2}Σ^{+}(v=0) state. Although inside a 2σ error estimate, these values are somewhat smaller than the experimentally observed value. In an effort to improve the precision of |μ|, the Stark effect in the B^{2}Σ^{+} - X^{2}Σ^{+} (0,0) band system of CaD has been recorded and analyzed. As expected, the Stark effect is about four times larger in CaD compared to CaH. The analysis gives |μ| of 2.633(27) D and 2.018(33) D for the X^{2}Σ^{+}(v=0) and B^{2}Σ^{+}(v=0) state, respectively. A discussion of the isotopic dependence of μ and experimental systematic errors will be given. The trend in μ for the X^{2}Σ^{+}, A^{2}Π and B^{2}Σ^{+} is modeled using a simple molecular orbital picture.

aSupported by NSF-Exp.Phys

bT. C. Steimle, J. Chen, and J. Gengler, J. Chem. Phys. 121, 829-834(2004).

cF. Holka and M. Urban, Chem. Phys. Lett. 426, 252(2006).

dI. S. K. Kerkines and A. Mavridis, J. Phys. Chem. A111, 371(2007).